OpenStudy (anonymous):

Width of rectangle is 8 more the length. The perimeter is 56. Help me solve with steps?

OpenStudy (anonymous):

Width steps LoL Word Play

OpenStudy (angela210793):

x=length y=width y=x+8 P=2x+2y=2x+2(x+8) 4x+16=56 4x=40 x=10

OpenStudy (anonymous):

ok now Find the dimentions >.< and then find the area?

OpenStudy (anonymous):

supposed tht length is x and width is x+8 2x+2(x+8)=56 4x+16=56 4x=40 x=10

OpenStudy (anonymous):

Hmm, usually the length is longer than the width. Still, if length is 10, then width is 18 and area is 180.